DEEP-UV CAVITY RING DOWN SPECTROSCOPY

M. SNEEP, S. HANNEMANN, E. J. VAN DUIJN, \underline{W. UBACHS}, Laser Centre, Department of Physics and Astronomy, Vrije Universiteit, Amsterdam, The Netherlands.

The sensitive optical detection technique of cavity ring-down spectroscopy is extended to wavelengths in the range 197– 204 nm, hence into the domain of the deep ultraviolet. A novel design narrow-band Fourier transform limited laser delivering laser pulses of 15 ns duration and a bandwidth of 50 MHz in the near-infrared is constructed, based on a passive Ti:Sa pulsed oscillator injection-seeded by a continuous-wave Ti:sa laser. Three stages of type-I frequency mixing produce effectively tunable radiation naar 200 nm at pulse energies of 0.1 mJ. The system is tested in a generic pulsed cavity-ring down setup to measure quantitatively the Rayleigh scattering cross section in SF_{6} and broadband extinction in CO_{2}. Resonant absorption was measured in the atmospheric Schumann-Runge (0,0), (1,0) and (2,0) bands of both 16O_{2} and ^{18}O_{2}. Besides an accurate determination of the molecular constants of the excited states this work also result in determination of vibrational and spin-dependent predissociation widths.